OpenStudy (justuu):

Simplify using the distributive property. 8(y + 12) 8y + 12 20y 20 + y 8y + 96

OpenStudy (justuu):

the last?

Nnesha (nnesha):

multiply both terms in the parentheses by outside number example \[\huge\rm a(b+c)=a *b +a*c=ab+ac \]

OpenStudy (justuu):

so is it the last one?

OpenStudy (♪chibiterasu):

yeah

OpenStudy (justuu):

thanks

Nnesha (nnesha):

and please post a little work it is hard for us to tell is it A , B ,C or d right or not work+ answer good job